OpenStudy (anonymous):

Have two tanks of 100 liters each, with salt solution. One is at 9% and 4% other. How many liters at least must be removed from each container, so that upon mixing which is removed to obtain a 7% solution of salt concentration?

OpenStudy (whpalmer4):

It's a little unclear what you are mixing: the two tanks start out with 100 liters each, one being a 9% salt solution and the other a 4% salt solution. After removing an unknown quantity from each tank, are you mixing the removed quantities, or the remaining quantities?

OpenStudy (mos1635):

which is removed to

OpenStudy (whpalmer4):

And there's no specified volume of 7% solution to be created?

OpenStudy (mos1635):

V1=volume of sol 1 V2= volume of sol 2 V1+V2=vo;ume of solution 3 \[C _{1}=\frac{ \frac{ 9 }{ Mr } }{ 0.1 }\] \[C _{2}=\frac{ \frac{ 4 }{ Mr } }{ 0.1 }\]

OpenStudy (mos1635):

\[C _{3}=\frac{ \frac{ 7 }{ Mr } }{ 0.1 }\]

OpenStudy (mos1635):

is that enough to get you started??